OpenStudy (anonymous):

Find an equation of the line that satisfies the given conditions.Through (−6, 5);parallel to the line x = 9

OpenStudy (owlcoffee):

for the first use: \[x-x _{1}=m(x-x _{1})\] We know that parallel lines have same slope, so you have the initial x, solve for Y.

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

That is too confusing

OpenStudy (owlcoffee):

Are you 100% that's the text of the problem?

OpenStudy (anonymous):

positive

OpenStudy (ybarrap):

x=9 is a vertical line. Any other vertical line will be parallel to this line. If you want the vertical line to go through point, (-6,5), then x=-6 would be a vertical line that also passes through the point (-6,5)

OpenStudy (anonymous):

so how would you get the equation?

OpenStudy (ybarrap):

The equation is simple: x=-6. This is an equation that includes all values of y, positive, negative and zero. It must. It is required to be parallel to x=9.

OpenStudy (ybarrap):

Maybe this makes the situation more clear: 

OpenStudy (ybarrap):

Does this make sense?

OpenStudy (anonymous):

Can you help me with another?

OpenStudy (anonymous):

Find an equation of the line that satisfies the given conditions. Through (−1, −2); perpendicular to the line 2x + 5y + 5 = 0

OpenStudy (owlcoffee):

Solved that one... http://openstudy.com/users/hannahmarie620#/updates/52338f82e4b0af32a078d43c